A boy is inside a bus moving at 5.0m/s toward the east. He is walking at 1.0 m/s toward the front. What is the velocity of the boy as observed by (a) a passenger sitting inside the bus and (b) another passenger waiting for the bua at a nearby bus stop?
1
Expert's answer
2019-08-23T11:11:56-0400

Velocity of boy towards east(u)="+5 \hat{i} (due \space to \space bus)+\hat{i}(due \space to \space his \space movement)=6 \hat{i}"

Velocity of fellow passenger(v)="+5 \hat{i} (due \space to \space bus)"

Velocity of waiting passenger(s)=0

(a):velocity of the boy as observed by fellow passenger sitting inside the bus=

u-v=

"6 \hat{i} -5 \hat{i}=\hat{i}"

That is 1 m/sec towards east.

(b) :velocity of the boy as observed by another passenger waiting for the bus at a nearby bus stop=

u-s = "6 \hat{i}-0=6\hat{i}"

Which means 6 m/sec towards east.